# Sharpness-Aware Minimization Can Hallucinate Minimizers

### v1

@@ Appendix B Omitted details for Section 3
B.1 Proof of Theorem 3.1
In this subsection, we prove Theorem 3.1. The argument relies on the implicit function theorem, which we state below.
Theorem B.1 (Implicit function theorem, Lee (2013)).
Let $U\subseteq\mathbb{R}^{d}\times\mathbb{R}^{d}$ be an open set, and let $(x,y)$ denote the coordinates on $U$. Suppose $\Phi:U\to\mathbb{R}^{d}$ is a smooth function, $(a,b)\in U$, and $c=\Phi(a,b)$. If the $d\times d$ matrix
$\left(\frac{\partial\Phi^{i}}{\partial y^{j}}(a,b)\right)$
is invertible, then there exist neighborhoods $V_{0}\subseteq\mathbb{R}^{d}$ of $a$ and $W_{0}\subseteq\mathbb{R}^{d}$ of $b$, together with a smooth function $F:V_{0}\to W_{0}$, such that $\Phi^{-1}(c)\cap(V_{0}\times W_{0})$ is the graph of $F$. In other words, $\Phi(x,y)=c$ for $(x,y)\in V_{0}\times W_{0}$ if and only if $y=F(x)$.
Now we prove Theorem 3.1.
@@
Let $x_{h}$ be a hallucinated minimizer of $f$ constructed in the proof of Theorem 2.2. Then, it satisfies
for some perturbation radius $\rho>0$. Let $V$ be an open neighborhood of $x_{h}$ on which the gradient never vanishes. Such a neighborhood exists because $\|\nabla f(x_{h})\|>0$ and $\nabla f$ is continuous. Define $F:\mathbb{R}^{d}\times V\rightarrow\mathbb{R}^{d}$ by
$F(x,y)=y+\rho\ u(y)-x.$
Clearly, $F(x_{\star},x_{h})=0$, and by assumption, $\frac{dF}{dy}(x_{\star},x_{h})=I+\rho\nabla u(x_{h})$ is invertible. By the implicit function theorem, there exist open neighborhoods $U_{0}\subseteq\mathbb{R}^{d}$ of $x_{\star}$ and $V_{0}\subseteq V$ of $x_{h}$, together with a smooth map $G:U_{0}\rightarrow V_{0}$, such that $G(x_{\star})=x_{h}$ and
$F(x,G(x))=0\quad\forall x\in U_{0}.$
This also implies that $G$ is also a local diffeomorphism at $x_{\star}$, since the differential is invertible at $x_{\star}$:
$I+\frac{dF}{dy}\bigg\rvert_{y=x_{h}}\frac{dG}{dx}\bigg\rvert_{x=x_{\star}}=0\iff\frac{dG}{dx}\bigg\rvert_{x=x_{\star}}=\left(\frac{dF}{dy}\right)\bigg\rvert_{y=x_{h}}^{-1}.$
In particular, there exists an open $U\subseteq U_{0}$ around $x_{\star}$ such that $G|_{U}:U\rightarrow G(U)$ is a diffeomorphism. Since $U\cap\mathcal{M}$ is an $m$-dimensional manifold (without boundary), the image $G(U\cap\mathcal{M})$ under the diffeomorphism is also an $m$-dimensional manifold.
Finally, note that any $y\in G(U\cap\mathcal{M})$ satisfies
$x=y+\rho\frac{\nabla f(y)}{\|\nabla f(y)\|},$
where $x=G^{-1}(y)\in\mathcal{M}\cap U$ and hence $x\in\operatorname*{arg\,min}{f}$. Thus, $G(U\cap\mathcal{M})$ forms an $m$-dimensional manifold of hallucinated minimizers. ∎
B.2 Proof of Theorem 3.2 and further discussion
See 3.2
Proof.
First, we claim that the set $H$ is closed (hence compact) by construction. Indeed, if $x_{h}\in\bar{H}$, then the corresponding function value must equal $f^{\mathrm{SAM}}(H)=\min f^{\mathrm{SAM}}$. By our assumption on $H$, this implies $x_{h}\in H$. Hence $H=\bar{H}$.
Let $\mathcal{N}_{\delta}(H)$ denote the closed $\delta$-neighborhood of $H$ from the theorem assumption. Since $1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x_{h})))>0$ and $\|\nabla f(x_{h})\|>0$ for all $x_{h}\in H$, there exists an open neighborhood $W$ of $H$ such that $1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x)))>0$ and $\|\nabla f(x)\|>0$ for any $x\in W$. By shrinking $\mathcal{N}_{\delta}(H)$ if necessary, we may assume $\mathcal{N}_{\delta}(H)\subseteq W$ and $f^{\mathrm{SAM}}$ is real-analytic on $\mathcal{N}_{\delta}(H)$.
@@
Then, $C_{\rho}$ is also compact. Set
$M:=\max_{x\in C_{\rho}}\|\nabla f(x)\|>0,\qquad L:=\max_{x\in\mathcal{N}_{\delta}(H)}\|\nabla^{2}f(x)\|>0,$
and
$\gamma:=\min_{x\in\mathcal{N}_{\delta}(H)}(1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x)))>0.$
Consider the SAM update with fixed $\rho>0$ and constant step size $\eta_{k}=\eta$ chosen such that
@@
$\nabla f^{\mathrm{SAM}}(x)=(I+\rho\nabla u(x))\nabla f(x^{+}),$
@@
$\leq f^{\mathrm{SAM}}(x_{k})-\eta\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x_{k})))\|\nabla f(x_{k}^{+})\|^{2}-\eta\|\nabla f(x_{k}^{+})\|^{2}+\frac{L\eta^{2}}{2}\|\nabla f(x_{k}^{+})\|^{2}$
$\leq f^{\mathrm{SAM}}(x_{k})-\eta\left(1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x_{k})))-\frac{L\eta}{2}\right)\|\nabla f(x_{k}^{+})\|^{2}$
@@
where the last inequality follows from $0<\eta<\frac{2\gamma}{L}$. Moreover, since the descent property $f^{\mathrm{SAM}}(x_{k+1})<f^{\mathrm{SAM}}(x_{k})$ holds when $\eta$ is replace by $\eta t$ for $t\in[0,1]$, the line segment from $x_{k}$ to $x_{k+1}$ lies within the sublevel set $(f^{\mathrm{SAM}})^{-1}([f_{\star},f_{\star}+\varepsilon])$. Thus, because $x_{k}\in C_{\varepsilon}$, we conclude $x_{k+1}\in C_{\varepsilon}$ by the connectedness of $C_{\varepsilon}$. This completes the induction.
Finally, since $f^{SAM}(x_{k})$ is decreasing and bounded below, we have $\eta\left(\gamma-\frac{L\eta}{2}\right)\|\nabla f(x_{k}^{+})\|\rightarrow 0$. Hence, $\nabla f(x_{k}^{+})\to 0$. If $x_{\infty}$ is a limit point of $\{x_{k}\}_{k=0,1,\dots}$, then $\nabla f(x_{\infty}^{+})=0$. By construction of $C_{\varepsilon}$, this implies $x_{\infty}\in H$. Therefore, $d(x_{k},H)\to 0$. ∎
@@
$1+\rho\lambda_{\mathrm{min}}(\mathrm{Sym}(\nabla u(x_{h})))>0.$
@@
Appendix C Experimental details for SAM in deep learning
C.1 SAM with full-batch gradients
Figure 7(a) presents a one-dimensional view along the line connecting $x_{h}$ and $x_{h}^{+}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}$, parameterized as $x(\alpha)=(1-\alpha)x_{h}+\alpha x_{h}^{+}$. The plot shows that $x_{h}$ is not a minimizer of the original objective and that the surrounding loss landscape differs substantially from that around $x_{h}^{+}$. This demonstrates that the phenomenon of SAM converging to a hallucinated minimizer is fundamentally distinct from the case in which a saddle point becomes an attractor, which requires the surrounding quadratic structure to hold (Compagnoni et al., 2023).
@@ C.1 SAM with full-batch gradients
Figure 7: Visualizations of the hallucinated minimizer $x_{h}$: (a) original loss $f$ along the line between $x_{h}$ and $x_{h}^{+}$; (b) SAM loss $f^{\mathrm{SAM}}$ over the plane defined by $x_{h}$, $x_{N}$, and $x_{N}^{\prime}$.
Figure 7(b) extends Figure 4(c) with an additional visualization of the SAM loss. We initialize $x_{0}$ by adding a small random perturbation of magnitude $0.1$ to $x_{h}$, and then perform $N=1000$ SAM steps, yielding the same $x_{N}$ reported in Figure 4(c). Applying an independent perturbation followed by the same procedure gives $x_{N}^{\prime}$. We then consider the plane spanned by $x_{h}$, $x_{N}$, and $x_{N}^{\prime}$, parameterized as $x(\alpha,\beta)=x_{h}+\alpha u+\beta v$ with $u=x_{N}-x_{h}$ and $v$ chosen orthogonal to $u$. On this plane, the visualization shows that the hallucinated minimizers are not confined to a one-dimensional curve but instead extend into a two-dimensional surface-like structure.
@@
Figure 9: Training loss and test accuracy curves for SAM with stochastic gradients on CIFAR-100 using ResNet-18. The top row shows SAM-only training, while the bottom row applies a switching strategy that runs stochastic gradient descent for the first 10% of epochs before switching to SAM. Columns correspond to perturbation radii $\rho\in\{0.1,0.4,0.7,1.0\}$. Final average test accuracies are 74.29%, 76.36%, 69.64%, and 61.58% for SAM-only, compared to 74.48%, 77.22%, 73.35%, and 67.93% for the switching strategy.
C.2 SAM with stochastic gradients
We examine whether the phenomena observed with full-batch SAM also arise in the stochastic setting, as shown in Figure 6. ResNet-18 is trained on CIFAR-100 with standard data augmentations, including random cropping with padding, horizontal flipping, and Cutout (DeVries & Taylor, 2017). The mini-batch size is $64$, the learning rate $0.01$, momentum $0.9$, and weight decay $10^{-4}$. Training proceeds for $200$ epochs with cosine-annealed learning rates (Loshchilov & Hutter, 2016), following the practical implementation of FSAM (Li et al., 2024). The switching strategy applies plain stochastic gradient descent for the first $10\%$ of epochs before switching to stochastic SAM.
Under the same setting, Figure 9 reports the CIFAR-100 results, comparing SAM-only with the switching strategy. Experiments are conducted for perturbation radii $\rho=0.1,0.4,0.7,1.0$ across $16$ random seeds. Each curve shows training loss and test accuracy over epochs, with bold lines denoting the mean across seeds and shaded regions indicating the standard deviation. The results show that, as in the full-batch case, SAM performance degrades with larger perturbation radii, whereas the switching strategy remains stable and robust across all settings.
Appendix D Two-dimensional synthetic function for visualization
To visualize how the SAM perturbation radius $\rho$ affects the objective, we introduce the following two-dimensional synthetic function (originally illustrated in Figure 1):
where
Figure 1 shows the original function $f$, the SAM objective $f^{\mathrm{SAM}}$, and its gradient $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$ at perturbation radius $\rho=2.8$. To further examine the effect of the perturbation radius, Figures 10 and 11 illustrate how the SAM objective $f^{\mathrm{SAM}}$ and its gradient $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$ evolve as $\rho$ varies over the range $0,0.5,\ldots,3.5$. In this setting, the SAM minimizers are defined as the regions where $f^{\mathrm{SAM}}$ attains its minimum values; in practice, they appear either as isolated points or as continuous curve-like structures.
@@ Appendix D Two-dimensional synthetic function for visualization
Figure 10: SAM objective $f^{\mathrm{SAM}}(x)$ under different perturbation radii $\rho$. The corresponding SAM minimizers are shown in pink.
@@
Figure 11: SAM gradient field $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$ for different perturbation radii $\rho$. The corresponding SAM minimizers are shown in pink.
An analysis of the SAM minimizers as a function of the perturbation radius $\rho$ reveals two distinct regimes. For small $\rho$, the minimizers approach the critical points of $f$. Although $f^{\mathrm{SAM}}$ is not defined at a critical point, higher-resolution numerical experiments show convergence arbitrarily close to such points. At $\rho=0$, $f^{\mathrm{SAM}}$ coincides with $f$, and the minimizers exactly match those of $f$. For $\rho=0.5$, the minimizers remain on the original minimizer set, whereas for $\rho=1.0$ and $\rho=1.5$, they shift toward the maximizers of $f$. The corresponding gradient fields indicate that these critical points act as attractors of the SAM dynamics, consistent with the theoretical analysis of Compagnoni et al. (2023).
In contrast, for larger perturbation radii, the SAM minimizers form a curve on the right-hand side, starting near the maximizer and drifting outward as $\rho$ increases. This behavior is consistent with the proof of Theorem 2.1, which shows that hallucinated minimizers emerge for a sufficiently large perturbation radius $\rho$. Theorem 3.1 further establishes that these minimizers preserve the dimensionality of the original minimizer manifold. Meanwhile, the SAM gradient field shows that these minimizers act as attractors within their neighborhood, and the conditions of Theorem 3.2 are indeed satisfied. Taken together, these observations show that our theory offers a full theoretical explanation of the empirical phenomena of hallucinated minimizers in this example.

### v2

@@ Appendix B Omitted details for Section 3
B.1 Proof of Theorem 3.1 and further discussion
@@
First, we claim that the set $H$ is closed (hence compact). Indeed, if $x_{h}\in\bar{H}$, then by continuity of $f^{\mathrm{SAM}}$ we have $f^{\mathrm{SAM}}(x_{h})=f^{\mathrm{SAM}}(H)=\min f^{\mathrm{SAM}}$. By the isolation assumption on the minimizer component, this implies $x_{h}\in H$. Hence $H=\bar{H}$.
Let $\mathcal{N}_{\delta}(H)$ denote the closed $\delta$-neighborhood of $H$ from the theorem assumption. Since
$1+\rho\lambda_{\min}(\mathrm{Sym}(\nabla u(x_{h})))>0$
and $\|\nabla f(x_{h})\|>0$ for all $x_{h}\in H$, there exists an open neighborhood $W$ of $H$ such that
$1+\rho\lambda_{\min}(\mathrm{Sym}(\nabla u(x)))>0$
and $\|\nabla f(x)\|>0$ for any $x\in W$. By shrinking $\mathcal{N}_{\delta}(H)$ if necessary, we may assume $\mathcal{N}_{\delta}(H)\subseteq W$ and $f^{\mathrm{SAM}}$ is real-analytic on $\mathcal{N}_{\delta}(H)$.
@@
Then, $C_{\rho}$ is also compact. Let $L>0$ be any constant such that $f^{\mathrm{SAM}}$ is $L$-smooth on $\mathcal{N}_{\delta}(H)$, and set $M:=\max_{x\in C_{\rho}}\|\nabla f(x)\|>0$. Let
$\gamma:=\min_{x\in\mathcal{N}_{\delta}(H)}\Bigl(1+\rho\,\lambda_{\min}(\mathrm{Sym}(\nabla u(x)))\Bigr).$
By continuity and compactness, $\gamma>0$. Consider the SAM update with fixed $\rho>0$ and constant step size $\eta_{k}=\eta$ chosen such that
@@
$\nabla f^{\mathrm{SAM}}(x)=(I+\rho\nabla u(x))^{\top}\nabla f(x^{+}),$
@@
$\leq f^{\mathrm{SAM}}(x_{k})-\eta\rho\lambda_{\min}(\mathrm{Sym}(\nabla u(x_{k})))\|\nabla f(x_{k}^{+})\|^{2}-\eta\|\nabla f(x_{k}^{+})\|^{2}+\frac{L\eta^{2}}{2}\|\nabla f(x_{k}^{+})\|^{2}$
$\leq f^{\mathrm{SAM}}(x_{k})-\eta\left(1+\rho\lambda_{\min}(\mathrm{Sym}(\nabla u(x_{k})))-\frac{L\eta}{2}\right)\|\nabla f(x_{k}^{+})\|^{2}$
@@
where the last inequality follows from $0<\eta<\frac{2\gamma}{L}$. Moreover, since the descent property $f^{\mathrm{SAM}}(x_{k+1})<f^{\mathrm{SAM}}(x_{k})$ holds when $\eta$ is replaced by $\eta t$ for $t\in[0,1]$, the line segment from $x_{k}$ to $x_{k+1}$ lies within the sublevel set $(f^{\mathrm{SAM}})^{-1}([f_{\star},f_{\star}+\varepsilon])$. Thus, because $x_{k}\in C_{\varepsilon}$, we conclude $x_{k+1}\in C_{\varepsilon}$ by the connectedness of $C_{\varepsilon}$. This completes the induction.
Finally, since $f^{\mathrm{SAM}}(x_{k})$ is decreasing and bounded below, we have $\eta\left(\gamma-\frac{L\eta}{2}\right)\|\nabla f(x_{k}^{+})\|^{2}\rightarrow 0$. Hence, $\nabla f(x_{k}^{+})\to 0$. If $x_{\infty}$ is a limit point of $\{x_{k}\}_{k=0,1,\dots}$, then $\nabla f(x_{\infty}^{+})=0$. By construction of $C_{\varepsilon}$, this implies $x_{\infty}\in H$. Therefore, $d(x_{k},H)\to 0$. ∎
@@
$1+\rho\lambda_{\min}(\mathrm{Sym}(\nabla u(x_{h})))>0.$
@@
B.2 Proof of Theorem 3.2
See 3.2
Proof.
Fix $\rho>0$ and define $\Phi_{\rho}(x):=x+\rho u(x)$ on the open set
$\Omega:=\{x\in\mathbb{R}^{d}:\ \nabla f(x)\neq 0\}.$
Since $f\in C^{2}$, the map $u(x)=\nabla f(x)/\|\nabla f(x)\|$ is $C^{1}$ on $\Omega$, and hence $\Phi_{\rho}$ is $C^{1}$ on $\Omega$ with Jacobian
$D\Phi_{\rho}(x)=I+\rho\nabla u(x).$
By assumption, $x_{h}\in\Omega$ and $D\Phi_{\rho}(x_{h})=I+\rho\nabla u(x_{h})$ is nonsingular. Therefore, by the inverse function theorem, there exist neighborhoods $V$ of $x_{h}$ and $U$ of $x_{\star}=\Phi_{\rho}(x_{h})$ such that
$\Phi_{\rho}:V\to U$
is a $C^{1}$ diffeomorphism with inverse $G:=\Phi_{\rho}^{-1}:U\to V$.
Now set
$\mathcal{H}:=G(\mathcal{M}\cap U)=\{x\in V:\ \Phi_{\rho}(x)\in\mathcal{M}\cap U\}.$
Since $\mathcal{M}\cap U$ is an embedded $m$-dimensional manifold and $G$ is a diffeomorphism, $\mathcal{H}$ is an embedded $m$-dimensional manifold as well.
@@
Figure 6: Visualizations of the hallucinated minimizer $x_{h}$: (a) original loss $f$ along the line between $x_{h}$ and $x_{h}^{+}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}$; (b) $f^{\mathrm{SAM}}$ over the affine plane through $x_{h}$ spanned by the directions $(x_{N}-x_{h})$ and $(x_{N}^{\prime}-x_{h})$.
@@
It remains to show that every $x\in\mathcal{H}$ is a hallucinated minimizer. Take any $x\in\mathcal{H}$. By definition, $x^{+}=\Phi_{\rho}(x)\in\mathcal{M}\subseteq\operatorname*{arg\,min}f$, so $\nabla f(x^{+})=0$ and $f(x^{+})=\min f$. Moreover, since $x\in V\subseteq\Omega$, we have $\nabla f(x)\neq 0$. Finally, $f^{\mathrm{SAM}}(x)=f(x^{+})=\min f$, and hence $x$ is a (global, hence local) minimizer of $f^{\mathrm{SAM}}$ with $\nabla f(x)\neq 0$ and $\nabla f(x^{+})=0$. Thus, $x$ is a hallucinated minimizer. ∎
Appendix C Further details on experiments
C.1 SAM with full-batch gradients on MNIST
Figure 6(a) presents a one-dimensional view along the line connecting $x_{h}$ and $x_{h}^{+}=x_{h}+\rho\frac{\nabla f(x_{h})}{\|\nabla f(x_{h})\|}$, parameterized as $x(\alpha)=(1-\alpha)x_{h}+\alpha x_{h}^{+}$. The plot shows that $x_{h}$ is not a minimizer of the original loss and that the surrounding loss landscape differs substantially from that around $x_{h}^{+}$. This demonstrates that the phenomenon of SAM converging to a hallucinated minimizer is fundamentally distinct from the case in which a saddle point becomes an attractor, which requires the surrounding quadratic structure to hold (Compagnoni et al., 2023).
Figure 6(b) extends Figure 3(c) with an additional visualization of the SAM loss. We initialize $x_{0}$ by adding a small random perturbation of magnitude $0.1$ to $x_{h}$, and then perform $N=1000$ SAM steps, yielding the same $x_{N}$ reported in Figure 3(c). Applying an independent perturbation followed by the same procedure gives $x_{N}^{\prime}$. We then consider the plane spanned by $x_{h}$, $x_{N}$, and $x_{N}^{\prime}$, parameterized as $x(\alpha,\beta)=x_{h}+\alpha d_{1}+\beta d_{2}$ with $d_{1}=x_{N}-x_{h}$ and $d_{2}$ chosen orthogonal to $d_{1}$. On this plane, the visualization shows that the hallucinated minimizers are not confined to a one-dimensional curve but instead extend into a two-dimensional surface-like structure.
@@ C.1 SAM with full-batch gradients on MNIST
Figure 8: Full-batch SAM on CIFAR-10 with ResNet-18: training loss $f(x)$ (left) and SAM loss $f^{\mathrm{SAM}}(x)$ (right) across perturbation radii $\rho$.
Figure 9: Full-batch SAM on CIFAR-100 with ResNet-18: training loss $f(x)$ (left) and SAM loss $f^{\mathrm{SAM}}(x)$ (right) across perturbation radii $\rho$.
@@
Figure 10: Training loss and test accuracy curves for SAM with stochastic gradients on CIFAR-100 using ResNet-18. Thin curves show individual seeds, and the bold curve shows the mean across seeds. The top row shows SAM-only training, while the bottom row applies a switching strategy that runs stochastic gradient descent for the first 10% of steps before switching to SAM. Columns correspond to perturbation radii $\rho\in\{0.1,0.4,0.7,1.0\}$. Final average test accuracies are 74.29%, 76.36%, 69.64%, and 61.58% for SAM-only, compared to 74.48%, 77.22%, 73.35%, and 67.93% for the switching strategy.
C.2 SAM with full-batch gradients on CIFAR-10/100
This subsection reports additional trajectories for the full-batch CIFAR-10/100 experiments in Section 4.1. We use full-batch gradients to isolate the shifted-gradient mechanism studied in our deterministic theory.
We train ResNet-18 from scratch on CIFAR-10/100 using full-batch SAM. To compute a full-batch gradient in practice, we accumulate gradients over mini-batches across the entire training set. The base optimizer is SGD with learning rate $10^{-3}$ and momentum $0.9$. We use standard per-channel normalization and no data augmentation to reduce confounding effects. We run 10,00010{,}000 SAM steps and sweep $\rho\in\{0.0,0.5,1.0,1.5,2.0\}$.
Figures 89 plot the trajectories of the training loss $f(x_{k})$ and the SAM loss $f^{\mathrm{SAM}}(x_{k})=f(x_{k}^{+})$. Across larger $\rho$, $f(x_{k})$ often plateaus above its smallest observed values, while $f(x_{k}^{+})$ remains much smaller, consistent with the mismatch signatures in Section 4.1.
C.3 SAM with stochastic gradients on CIFAR-100
We examine how SAM behaves under standard mini-batch training on CIFAR-100 as the perturbation radius $\rho$ varies, as summarized in Figure 5. ResNet-18 is trained with standard data augmentations, including random cropping with padding, horizontal flipping, and Cutout (DeVries and Taylor, 2017). The mini-batch size is $64$, the learning rate $0.01$, momentum $0.9$, and weight decay $10^{-4}$. Training proceeds for $200$ epochs with cosine-annealed learning rate, following the implementation details in Li et al. (2024). We compare SAM-only against a short SGD warm-start before enabling SAM: we run plain SGD for the first $10\%$ of epochs and then continue with (stochastic) SAM for the remainder.
Under this setting, Figure 10 reports representative loss/accuracy curves at $\rho\in\{0.1,0.4,0.7,1.0\}$ over $16$ random seeds. Each curve shows training loss and test accuracy over epochs, with bold lines denoting the mean across seeds. Overall, as $\rho$ increases, SAM-only training becomes less stable and more sensitive to the choice of $\rho$, whereas the SGD warm-start improves robustness across the same sweep.
@@ C.3 SAM with stochastic gradients on CIFAR-100
Figure 11: SAM loss $f^{\mathrm{SAM}}(x)$ under different perturbation radii $\rho$. The corresponding minimizers of $f^{\mathrm{SAM}}$ are shown in pink.
@@
Figure 12: SAM gradient field $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$ for different perturbation radii $\rho$. The corresponding minimizers of $f^{\mathrm{SAM}}$ are shown in pink.
Appendix D Two-dimensional synthetic function for visualization
To visualize how the SAM perturbation radius $\rho$ affects the SAM loss $f^{\mathrm{SAM}}$, we introduce the following two-dimensional synthetic function (originally illustrated in Figure 1):
where
Figure 1 shows the original function $f$, the SAM loss $f^{\mathrm{SAM}}$, and the SAM gradient $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$ at perturbation radius $\rho=2.8$. To further examine the effect of the perturbation radius, Figures 11 and 12 illustrate how the SAM loss $f^{\mathrm{SAM}}$ and the SAM gradient $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$ evolve as $\rho$ varies over the range $0,0.5,\ldots,3.5$. In this setting, the minimizers of $f^{\mathrm{SAM}}$ are defined as the regions where $f^{\mathrm{SAM}}$ attains its minimum values; in practice, they appear either as isolated points or as continuous curve-like structures.
An analysis of the minimizers of $f^{\mathrm{SAM}}$ as a function of the perturbation radius $\rho$ reveals two distinct regimes. For small $\rho$, the minimizers approach the critical points of $f$. Although $f^{\mathrm{SAM}}$ is not defined at a critical point, higher-resolution numerical experiments show convergence arbitrarily close to such points. At $\rho=0$, $f^{\mathrm{SAM}}$ coincides with $f$, and the minimizers exactly match those of $f$. For $\rho=0.5$, the minimizers remain on the original minimizer set, whereas for $\rho=1.0$ and $\rho=1.5$, they shift toward the maximizers of $f$. The corresponding gradient fields indicate that these critical points act as attractors of the SAM dynamics, consistent with the theoretical analysis of Compagnoni et al. (2023).
In contrast, for larger perturbation radii, the minimizers of $f^{\mathrm{SAM}}$ form a curve on the right-hand side, starting near the maximizer and drifting outward as $\rho$ increases. This behavior is consistent with the proof of Theorem 2.2, which shows that hallucinated minimizers emerge for a large $\rho$. Theorem 3.2 further establishes that these minimizers preserve the dimensionality of the original minimizer manifold. Meanwhile, the SAM gradient field shows that these minimizers act as attractors within their neighborhood, and the conditions of Theorem 3.1 are indeed satisfied.
Taken together, this 2D example illustrates three qualitative regimes: (i) for small $\rho$, minimizers of $f^{\mathrm{SAM}}$ track minimizers/critical regions of $f$ (with the caveat that $f^{\mathrm{SAM}}$ is undefined exactly at critical points); (ii) as $\rho$ increases, the minimizers of $f^{\mathrm{SAM}}$ can drift toward high-loss regions near maximizers, consistent with the large-$\rho$ mechanism in Theorem 2.2; and (iii) the resulting minimizers of $f^{\mathrm{SAM}}$ can form non-isolated, curve-like sets, consistent with the manifold mechanism in Theorem 3.2. Moreover, the SAM gradient field visually points toward these minimizer sets, suggesting local attraction in the sense of Theorem 3.1 (the eigenvalue condition can be checked numerically in this low-dimensional setting).